# New Convergence Analysis of GMRES with

### v1

@@ 6 Hpd preconditioning for $\mathbf{A}$ positive definite
In the proof of Theorem 6, in each case, the two matrices ${\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\pgfsys@color@rgb@stroke{0}{0.65}{0}\pgfsys@color@rgb@fill{0}{0.65}{0}\mathbf{M}^{-1}}$ can be replaced by any hpd matrix, and the thesis of the theorem changed appropriately.
@@ 7.1 Choice of deflation space and convergence of WPD-GMRES
Given a pd matrix $\mathbf{A}$, let $(\lambda_{j},\mathbf{z}^{(j)})_{j=1,\dots n}$ denote the eigenpairs of generalized eigenvalue problem (4), i.e., $\mathbf{N}\mathbf{z}^{(j)}=\lambda_{j}\mathbf{M}\mathbf{z}^{(j)}$, with $\mathbf{M}$ and $\mathbf{N}$ the Hermitian and skew-Hermitian parts of $\mathbf{A}$ as in (5). Let $\tau>0$. Consider the space
@@
Let $\mathcal{Z}$ be as in Definition 2. Its orthogonal complement is the space $\mathcal{Z}^{\perp}=\{\mathbf{M}\mathbf{z}^{(k)};|\lambda_{k}|\leq\tau\}.$ Moreover, any $\mathbf{y}\in\mathcal{Z}^{\perp}\setminus\{\mathbf{0}\}$ satisfies
@@
The case where no vectors are deflated corresponds to setting $\tau=\rho(\mathbf{M}^{-1}\mathbf{N})$ (the spectral radius). In that case, the estimate in Theorem 6 is exactly the estimate in [22, Corollary 4.4].
@@ 7.2 Real-valued Case
8 Numerical Illustration: Convection-Diffusion-Reaction
In this section, the problem considered is the convection-diffusion-reaction problem posed in $\Omega=[-1,1]^{2}$. It is a real-valued problem ($\mathbb{K}=\mathbb{R}$), so Hermitian means symmetric. The strong formulation of the problem is:
@@ 8 Numerical Illustration: Convection-Diffusion-Reaction
It can be remarked that $\operatorname{div}\mathbf{a}=0$.
Figure 1: Example meshes. Left: 2373 vertices and 4568 triangles. Right: 8643 vertices and 16948 triangles.
Figure 2: Left: Solution. Right: Partition into 16 subdomains for $\mathbf{H}_{\text{DD}}$
@@
The choice of $\mathbf{H}={{\mathbf{M}}^{-1}}$ was used, e.g., in [1]. It is also a fundamental feature of the CGW method [7, 26, 28], and was successfully used recently for the solution of Port-Hamiltonian systems [13].
@@
where $k_{0}$ denotes the maximal number of subdomains that each mesh element belongs to [21, Theorem 4.40] and $\tau^{\prime}$ is a parameter that has been set to $0.15$. The constant in the bound does not depend on the total number $N$ of subdomains or the mesh parameter $h$.
@@
Deflation Operator.
We aim to illustrate the convergence result in Theorem 6. Once $\mathbf{M}$ and $\mathbf{N}$ have been assembled by FreeFem++, they are imported into Octave. The generalized eigenvalue problem (4) (i.e., $\mathbf{N}\mathbf{z}^{(j)}=\lambda_{j}\mathbf{M}\mathbf{z}^{(j)}$) is partially solved by eigs: the eigenpairs corresponding to the eigenvalues of largest magnitude are approximated. The eigenpairs are ordered in decreasing order of magnitude of $\lambda_{j}$. Since our problem is real-valued, the eigenvectors can be grouped into complex conjugate pairs and we apply the strategy from Theorem 7 to obtain real-valued deflation operators. The Octave command is $\mathbf{Z}$ = [real($\mathbf{V}$(:,1:2:m)), imag($\mathbf{V}$(:,1:2:m))], where the columns of $\mathbf{V}$ are the sorted eigenvectors and m is the (even) number of vectors that should form the deflation space. Finally, we set $\mathbf{Y}=\mathbf{H}\mathbf{A}\mathbf{Z}$ and compute the deflation operators as in Definition 1.
@@
[…]
Quantities of interest.
Figure 3: $|\lambda_{1}|$ to $|\lambda_{3000}|$ in log scale (solution of (4))
@@
Table 1: If m is the rank of the deflation space, the second line is the modulus of the first eigenvalue not included in the deflation space, and the next three lines correspond to the theoretical bound $\theta_{th}$ for out three choices of preconditioner.
Results for $\eta=1$.
For this test case, we use a mesh with 63658 triangles and 32158 vertices. Once the homogeneous Dirichlet boundary condition has been treated by elimination, the problem has 31502 degrees of freedom. The parameter in the convection field is $\eta=1$ so that $\mathbf{A}=\mathbf{M}+\mathbf{N}$, with $\mathbf{N}=\tilde{}\mathbf{N}$. We first solve generalized eigenproblem (4). The spectrum is represented in Figure 3 where the magnitudes of the 3000 largest (purely imaginary) eigenvalues are represented. It is slightly disappointing that the distribution of eigenvalues does not consist in a cluster of large eigenvalues and a tail of eigenvalues tightly clustered around 0. This would indeed have been ideal for selecting a value of $\tau$ to compute the deflation space with the formula in Definition 2. Table 1 gives the value of $|\lambda_{\text{m}+1}|$ for a few choices of m as well as the resulting bounds for $\theta$. In these bounds we have injected the approximate condition number $\kappa(\mathbf{H}\mathbf{M})$ which takes the following values
@@
0 671 7.133e-05 1.658e-02
10 543 9.166e-05 2.998e-02
50 440 9.798e-05 4.152e-02
100 400 9.938e-05 4.772e-02
200 369 1.002e-04 5.229e-02
500 321 1.007e-04 6.847e-02
1000 282 1.009e-04 8.047e-02
Figure 4: Convergence curves when $\eta=1$ with varying preconditioner and rank of deflation space
Results for $\eta=100$.
Next we change the value of $\eta$ to $100$. The global matrix is now $\mathbf{A}=\mathbf{M}+\mathbf{N}$ with $\mathbf{N}=100\tilde{}\mathbf{N}$, and the problem is much more nonsymmetric. The eigenvalues arising from (4) get multiplied by $100$ which has the effect of seriously deteriorating the bounds for $\theta$. Figure 5 shows the convergence curves in this case. Again, we confirm that applying a preconditioner tailored for the symmetric part is a good idea (with 341 or 352 iterations instead of 1276). In the non-preconditioned case, it occurs that the non-deflated problem converges the fastest. This is surprising but does not contradict the theory. Figure 6 shows the first 1000 residuals in this case for m$=0$ and m$=100$. It can be observed that deflation initially accelerates convergence, particularly where it is slowest. However, after roughly 500 iterations, the non-deflated algorithm is faster. For $\mathbf{H}=\mathbf{H}_{\text{DD}}$ or ${{\mathbf{M}}^{-1}}$, convergence improves when more vectors are added to the deflation space. With $\mathbf{H}={{\mathbf{M}}^{-1}}$, deflating 10 vectors reduces the iteration count from 341 to 331 and deflating 100 vectors reduces it further to 259. With $\mathbf{H}=\mathbf{H}_{\text{DD}}$, deflating 10 vectors reduces the iteration count from 352 to 343 and deflating 100 vectors reduces it further to 275.
@@
0 1276 2.424e-08 4.447e-03
10 1740 9.858e-08 3.898e-03
50 1863 3.219e-07 4.036e-03
100 1835 5.980e-07 4.058e-03
200 1865 1.151e-06 5.914e-03
500 1688 2.825e-06 7.266e-03
1000 1470 5.828e-06 9.225e-03
Figure 5: Convergence curves when $\eta=100$ with varying preconditioner and rank of deflation space
Figure 6: Zoom on the first 1000 iterations in the case $\mathbf{H}=\mathbf{{I}}$ and $\eta=100$.
@@
We consider the case where $\eta=100$ and $\mathbf{H}=\mathbf{H}_{\text{DD}}$ with varying mesh size. The stopping criterion is as before, $\|\mathbf{r}_{i}\|_{\mathbf{H}}/\|\mathbf{r}_{0}\|_{\mathbf{H}}=\|\mathbf{r}_{i}\|_{\mathbf{H}}/\|\mathbf{b}\|_{\mathbf{H}}<10^{-10}$. The three considered meshes are the $32158$ vertex mesh that was used in the tests above as well as the two less refined meshes represented in Figure 1 (2373 and 8643 vertices). After elimination of the degrees of freedom on the boundary, the resulting linear systems have respectively 31502, 8307 and 2197 dofs. The problem is solved by WPD-GMRES without deflation ($\text{m}=0$ and with deflation of $\text{m}=100$ vectors). The results are presented in Table 2. We notice that the iteration counts increase weakly with the mesh size. For example, from 236 for 2,197 dofs to 275 for 32,502 dofs in the case where 100 vectors are deflated. We also report on the quantities that appear in the convergence bounds: $\kappa(\mathbf{H}_{\text{DD}}\mathbf{M})$, $|\lambda_{1}|=\rho(\mathbf{M}^{-1}\mathbf{N})$ and $|\lambda_{101}|$. We know from the theory in [24] that $\kappa(\mathbf{H}_{\text{DD}}\mathbf{M})$ is bounded independently of the mesh size and we indeed notice that it does not depend very much on the mesh. In [22, Section 5], it was proved that for this particular PDE, $|\lambda_{1}|=\rho(\mathbf{M}^{-1}\mathbf{N})$ is also bounded independently of the mesh size by
@@
Table 2: For $\eta=100$ and $\mathbf{H}=\mathbf{H}_{\text{DD}}$, three mesh sizes are considered
@@
Table 3: Iteration counts for weighted and unweighted algorithms. Both algorithms use the stopping criterion coming from the unweighted algorithm
Figure 7: Comparison of weighted and unweighted algorithms, without deflation and with deflation of $100$ vectors. The chosen norm is the one that is minimized by the unweighted algorithm

### v2

@@ 5 The case of hpd preconditioning for $\mathbf{A}$ positive definite
In the proof of Theorem 6, in each case, the two matrices ${\color[rgb]{0,0.65,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.65,0}\mathbf{M}^{-1}}$ can be replaced by any hpd matrix, say $\tilde{\mathbf{M}}^{-1}$, and the thesis of the theorem changed appropriately so that, e.g. for case 1,
$\theta(\mathbf{A},\mathbf{H},\mathbf{H},\mathbf{H}\mathbf{A}\mathbf{Z},\mathbf{Z})\geq\frac{\lambda_{\min}(\mathbf{H}\mathbf{M})}{\lambda_{\max}(\mathbf{H}\tilde{}\mathbf{M})}\times\operatorname{inf}\limits_{\mathbf{y}\in\ker(\mathbf{Z}^{*}\mathbf{A}^{*}\mathbf{A}^{-1})\setminus\{\mathbf{0}\}}\frac{|\langle\mathbf{A}^{-1}\mathbf{y},\mathbf{y}\rangle|}{\langle\mathbf{y},{\tilde{\mathbf{M}}^{-1}}\mathbf{y}\rangle}\cdot$
@@ 6.1 Choice of deflation space and convergence of WPD-GMRES
Given a pd matrix $\mathbf{A}$, let $(\lambda_{j},\mathbf{z}^{(j)})_{j=1,\dots n}$ denote the eigenpairs of the generalized eigenvalue problem (4), i.e., $\mathbf{N}\mathbf{z}^{(j)}=\lambda_{j}\mathbf{M}\mathbf{z}^{(j)}$, with $\mathbf{M}$ and $\mathbf{N}$ the Hermitian and skew-Hermitian parts of $\mathbf{A}$ as in (5). Let $\tau>0$. Consider the space
@@
From Lemma 1 it follows that this generalized eigenvalue problem has a complete set of eigenvactors, and this is used in the following result.
@@
Let $\mathcal{Z}$ be as in Definition 2. Its orthogonal complement is the space $\mathcal{Z}^{\perp}=\operatorname{span}\{\mathbf{M}\mathbf{z}^{(k)};|\lambda_{k}|\leq\tau\}.$ Moreover, any $\mathbf{y}\in\mathcal{Z}^{\perp}\setminus\{\mathbf{0}\}$ satisfies
@@
In the following we use the short-hand notation $\mathbf{A}^{-*}$ for $(\mathbf{A}^{*})^{-1}=(\mathbf{A}^{-1})^{*}$.
@@
The case where no vectors are deflated corresponds to setting $\tau=\rho(\mathbf{M}^{-1}\mathbf{N})$ (the spectral radius). In that case, the estimate in Theorem 6 is exactly the estimate in [26, Corollary 4.4].
@@ 6.2 Real-valued Case
7 Numerical experiments
We present three different types of numerical experiments illustrating how the proposed deflation space, inspired by our new bounds, can reduce the number of iterations needed for convergence. In order for the proposed deflation to be practical, either the cost of computing the eigenvectors of the generalized skew-Hermitian problem (4) is offset by the saving in number of iterations, or, as in our second example, when GMRES fails to converge, but with a relatively small deflation space, convergence is attained.
Since our problems are real-valued ($\mathbb{K}=\mathbb{R}$), Hermitian means symmetric. Moreover, the eigenvectors can be grouped into complex conjugate pairs and we apply the strategy from Theorem 7 to obtain real-valued deflation operators. The Octave command is $\mathbf{Z}$ = [real($\mathbf{V}$(:,1:2:m)), imag($\mathbf{V}$(:,1:2:m))], where the columns of $\mathbf{V}$ are the sorted eigenvectors and m is the (even) number of vectors that should form the deflation space. Finally, we set $\mathbf{Y}=\mathbf{H}\mathbf{A}\mathbf{Z}$ and compute the deflation operators as in Definition 1.
Quantities of interest.
7.1 Numerical Illustration: Jordan block
We first consider an academic problem. Namely, a $1000\times 1000$ (bidiagonal) scaled Jordan block of the form
$\mathbf{A}=\begin{pmatrix}1&\alpha&&&&\\ &1&\alpha&&&\\ &&\ddots&\ddots&&\\ &&&1&\alpha\\ &&&&1\end{pmatrix}$
for which the symmetric part $\mathbf{M}$ is a tridiagonal Toeplitz matrix. Choosing $\alpha=0.99$ guarantees that all eigenvalues of the symmetric part $\mathbf{M}$ are positive so $\mathbf{A}$ is positive definite; see, e.g., [19, 1.4.I16]. The condition number of the symmetric part is $\kappa(\mathbf{M})=199$. Our objective is to illustrate one of the main contribution of the article which is the new deflation space. For this reason we do not consider preconditioning or weighting, i.e., $\mathbf{W}=\mathbf{H}=\mathbf{{I}}$. Figure 1 shows the convergence curves obtained without deflation and with deflation of a varying number $m$ of vectors. This confirms that the vectors arising from the generalized eigenvalue problem (4) are good candidates for deflation. Deflating 50 eigenvectors decreases the iteration count by 348 iteration and deflating 100 eigenvectors decreases the iteration count by 600. In Table 1 we show for various values of $m$, the dimension of the deflation space, number of iterations to converge, the quantities $\theta_{exp}$ calculated from the experiment, and $\theta_{th}$ the bound for them as in Theorem 4.
Figure 1: Convergence for the scaled Jordan block example (Section 7.1). The number $m$ of deflated vectors varies. Increasing $m$ accelerates convergence as predicted by the theory.
$m$ iter $\theta_{th}$ $\theta_{exp}$
0 1000 1.00e-04 1.99e-02
10 959 1.02e-04 1.99e-02
50 652 1.33e-04 1.99e-02
100 400 2.25e-04 1.99e-02
200 188 5.79e-04 2.00e-02
300 110 1.13e-03 1.99e-02
400 73 1.81e-03 2.09e-02
500 51 2.58e-03 2.38e-02
Table 1: Scaled Jordan block example (Section 7.1). $m$ is the number of deflated vectors, iter is the iteration count, $\theta_{th}$ and $\theta_{exp}$ are the theoretical lower bound and observed worse value of $\theta$.
7.2 Numerical Illustration: Stokes problem
We consider the incompressible Stokes equation as in [15, Section 6.2]. The problem is discretized using the $\mathbb{Q}_{1}-\mathbb{Q}_{1}$ finite element approximation of the (unsteady) channel domain problem in IFISS [23]. The system matrices are of the form $\mathbf{A}=\mathbf{M}+\mathbf{N}$ with
$\mathbf{M}=\begin{pmatrix}\mathcal{M}-\tau/2\mathbf{A}_{H}&\mathbf{0}\\ \mathbf{0}&-\tau/2\mathbf{C}\end{pmatrix}\text{ and }\mathbf{N}=\begin{pmatrix}\mathbf{0}&-\tau/2\mathbf{B}\\ \tau/2\mathbf{B}^{*}&0\end{pmatrix}$
where $\tau/2=10^{-4}$ is the time step, $\mathcal{M}$ is the mass matrix, $\mathbf{A}_{H}$ is the discretization of the second order term in the PDE, $\mathbf{C}$ is a stabilization term and $\mathbf{B}$ is the discrete divergence. The total number of degrees of freedom is $n=12,675$ and the size of the $\mathbf{B}$ block is $8,450\times 4,225$. The convergence curves are presented in Figure 2 where it can be observed that deflation of 1000 vectors allows to converge to a relative tolerance of $10^{-10}$ in 877 iterations whereas the non deflated problem does not converge. Table 2 shows that the theoretical $\theta_{th}$ does not vary much with the size $m$ of the deflation space. The condition number of the Hermitian part is very large: $\kappa(\mathbf{M})=8.93e6$.
$m$ iter $\theta_{th}$ $\theta_{exp}$
0 1000 1.12e-07 2.07e-04
10 1000 1.12e-07 3.41e-03
50 1000 1.12e-07 8.43e-03
500 1000 1.12e-07 4.08e-03
1000 877 1.12e-07 2.28e-03
Table 2: Stokes example (Section 7.2). $m$ is the number of deflated vectors, iter is the iteration count, $\theta_{th}$ and $\theta_{exp}$ are the theoretical lower bound and observed worse value of $\theta$.
Figure 2: Convergence for the Stokes problem (Section 7.2). The number $m$ of deflated vectors varies. Increasing $m$ improves convergence as predicted by the theory.
7.3 Numerical Illustration: Convection-Diffusion-Reaction
In this section, the problem considered is the convection-diffusion-reaction problem posed in $\Omega=[-1,1]^{2}$. The strong formulation of the problem is:
@@ 7.3 Numerical Illustration: Convection-Diffusion-Reaction
It can be remarked that $\operatorname{div}\mathbf{a}=0$. This is the same PDE as in [26] with a change of variables.
Figure 3: Example meshes for the convection-diffusion-reaction problem (Section 7.3). Left: 2373 vertices and 4568 triangles. Right: 8643 vertices and 16948 triangles.
Figure 4: Convection-diffusion-reaction problem (Section 7.3). Left: Solution. Right: Partition into 16 subdomains for $\mathbf{H}_{\text{DD}}$.
@@
The choice of $\mathbf{H}={{\mathbf{M}}^{-1}}$ was used, e.g., in [1]. It is also a fundamental feature of the CGW method [7, 30, 32], and was successfully used recently for the solution of Port-Hamiltonian systems [15]. In our experiments we used a direct solver (Octave’s backslash) to apply $\mathbf{M}^{-1}$.
@@
where $k_{0}$ denotes the maximal number of subdomains that each mesh element belongs to [25, Theorem 4.40] and $\tau^{\prime}$ is a parameter that has been set to $0.15$. The constant in the bound does not depend on the total number $N$ of subdomains or the mesh parameter $h$.
@@
Deflation Operator
We aim to illustrate the convergence result in Theorem 6. Once $\mathbf{M}$ and $\mathbf{N}$ have been assembled by FreeFem++, they are imported into Octave. The generalized eigenvalue problem (4) (i.e., $\mathbf{N}\mathbf{z}^{(j)}=\lambda_{j}\mathbf{M}\mathbf{z}^{(j)}$) is partially solved by eigs: the eigenpairs corresponding to the eigenvalues of largest magnitude are approximated.
@@
When $\mathbf{H}=\mathbf{H}_{\text{DD}}$ is applied, the preconditioner already contains a deflation operator $\boldsymbol{\Pi}$ which corresponds to a deflation space formed by eigenvectors of frequency less than a chosen $\tau^{\prime}$ of well chosen eigenproblems in the subdomains. The presence of a deflation operator and deflation space in $\mathbf{H}_{\text{DD}}$ does not interfere with the computation of a deflation space and deflation operator following the definition in Definition 2.
Figure 5: Convection-diffusion-reaction problem (Section 7.3). $|\lambda_{1}|$ to $|\lambda_{3000}|$ in log scale (solution of (4) for $\eta=1$).
@@
Table 3: Convection-diffusion-reaction problem (Section 7.3). If m is the rank of the deflation space, the second line is the modulus of the first eigenvalue not included in the deflation space, and the next three lines correspond to the theoretical bound $\theta_{th}$ for out three choices of preconditioner when $\eta=1$.
Results for $\eta=1$
For this test case, we use a mesh with 63658 triangles and 32158 vertices. Once the homogeneous Dirichlet boundary condition has been treated by elimination, the problem has 31502 degrees of freedom. The parameter in the convection field is $\eta=1$ so that $\mathbf{A}=\mathbf{M}+\mathbf{N}$, with $\mathbf{N}=\tilde{}\mathbf{N}$. We first solve generalized eigenproblem (4). The spectrum is represented in Figure 5 where the magnitudes of the 3000 largest (purely imaginary) eigenvalues are represented. Unfortunately, the distribution of eigenvalues does not consist in a cluster of large eigenvalues and a tail of eigenvalues tightly clustered around 0. This would indeed have been ideal for selecting a value of $\tau$ to compute the deflation space with the formula in Definition 2. Table 3 gives the value of $|\lambda_{\text{m}+1}|$ for a few choices of m as well as the resulting bounds for $\theta_{th}$. In these bounds we have injected the approximate condition number $\kappa(\mathbf{H}\mathbf{M})$ which takes the following values
@@
0 671 7.1e-05 1.7e-02
10 543 9.2e-05 3.0e-02
50 440 9.8e-05 4.2e-02
100 400 9.9e-05 4.8e-02
200 369 1.0e-04 5.2e-02
500 321 1.0e-04 6.8e-02
1000 282 1.0e-04 8.0e-02
Figure 6: Convergence curves when $\eta=1$ in Section 7.3 with varying preconditioner and rank of deflation space.
Results for $\eta=100$
Next we change the value of $\eta$ to $100$. The global matrix is now $\mathbf{A}=\mathbf{M}+\mathbf{N}$ with $\mathbf{N}=100\tilde{}\mathbf{N}$, and the problem is much more nonsymmetric. The eigenvalues arising from (4) get multiplied by $100$ which has the effect of seriously deteriorating the bounds for $\theta_{th}$. Figure 7 shows the convergence curves in this case. Again, we confirm that applying a preconditioner tailored for the symmetric part is a good idea (with 341 or 352 iterations instead of 1276). In the non-preconditioned case, it occurs that the non-deflated problem converges the fastest. This is surprising but does not contradict the theory. Figure 8 shows the first 1000 residuals in this case for m$=0$ and m$=100$. It can be observed that deflation initially accelerates convergence, particularly where it is slowest. However, after roughly 500 iterations, the non-deflated algorithm is faster. For $\mathbf{H}=\mathbf{H}_{\text{DD}}$ or ${{\mathbf{M}}^{-1}}$, convergence improves when more vectors are added to the deflation space. With $\mathbf{H}={{\mathbf{M}}^{-1}}$, deflating 10 vectors reduces the iteration count from 341 to 331 and deflating 100 vectors reduces it further to 259. With $\mathbf{H}=\mathbf{H}_{\text{DD}}$, deflating 10 vectors reduces the iteration count from 352 to 343 and deflating 100 vectors reduces it further to 275.
@@
0 1276 2.4e-08 4.4e-03
10 1740 9.9e-08 3.9e-03
50 1863 3.2e-07 4.0e-03
100 1835 6.0e-07 4.1e-03
200 1865 1.2e-06 5.94e-03
500 1688 2.8e-06 7.3e-03
1000 1470 5.8e-06 9.2e-03
Figure 7: Convergence curves when $\eta=100$ in Section 7.3 with varying preconditioner and rank of deflation space.
Figure 8: Zoom on the first 1000 iterations in the case $\mathbf{H}=\mathbf{{I}}$ and $\eta=100$. (Section 7.3).
Comparison with the Elman bound
A very well known bound for GMRES is the so-called Elman bound [8, 9]. Its immediate generalization to GMRES preconditioned by $\mathbf{H}$ and weighted by $\mathbf{W}$ states that $\frac{\|\mathbf{r}_{i}\|_{\mathbf{W}}^{2}}{\|\mathbf{r}_{i-1}\|_{\mathbf{W}}^{2}}\leq 1-\theta_{\text{El}}(\mathbf{A},\mathbf{H},\mathbf{W})$ where
$\theta_{\text{El}}(\mathbf{A},\mathbf{H},\mathbf{W})=\left(\frac{\mathrm{d}(0,W_{\mathbf{W}}(\mathbf{A}\mathbf{H}))}{\|\mathbf{A}\mathbf{H}\|_{\mathbf{W}}}\right)^{2}.$
The numerator is the distance between 00 and the $\mathbf{W}$-field of value of the preconditioned operator. When $\mathbf{W}=\mathbf{H}$ for some hpd matrix $\mathbf{H}$, it can be shown that
$\theta_{\text{El}}(\mathbf{A},\mathbf{H},\mathbf{H})=\frac{\lambda_{\min}(\mathbf{H}\mathbf{M})^{2}}{\lambda_{\max}(\mathbf{A}^{*}\mathbf{H}\mathbf{A}\mathbf{H})},$
Where we can use the $\lambda_{\max}$ notation since the eigenvalues are real and positive. In Table 4, for each preconditioner and for $\eta\in\{1,100\}$ we compare $\theta_{th}$ and $\theta_{exp}$ to $\theta_{\text{El}}$. For the very particular case where $\mathbf{W}=\mathbf{H}=\mathbf{M}^{-1}$, it can be proved that both bounds are identical. For the other cases, it can be observed that our bound is lower than the Elman bound.
$\theta_{exp}$ $\theta_{th}$ $\theta_{\text{El}}$
$\mathbf{H}={{\mathbf{M}}^{-1}}$ $\eta=1$ 7.956e-01 7.059e-01 7.059e-01
$\eta=100$ 3.144e-03 2.399e-04 2.399e-04
$\mathbf{H}=\mathbf{H}_{\text{DD}}$ $\eta=1$ 5.314e-01 4.346e-02 2.399e-04
$\eta=100$ 7.453e-03 1.477e-05 9.321e-06
$\mathbf{H}=\mathbf{{I}}$ $\eta=1$ 1.658e-02 7.133e-05 1.021e-08
$\eta=100$ 4.447e-03 2.424e-08 7.774e-09
Table 4: For the convection-diffusion-reaction problem in Section 7.3 without deflation, comparison between $\theta_{exp}=\min_{i}(1-{\|\mathbf{r}_{i}\|_{\mathbf{W}}^{2}}/{\|\mathbf{r}_{i-1}\|_{\mathbf{W}}^{2}}$), $\theta_{th}$ the bound from this paper (also the bound in [26]) and $\theta_{\text{El}}$, the Elman bound. Larger $\theta$ means faster convergence.
@@
We consider the case where $\eta=100$ and $\mathbf{H}=\mathbf{H}_{\text{DD}}$ with varying mesh size. The stopping criterion is as before, $\|\mathbf{r}_{i}\|_{\mathbf{H}}/\|\mathbf{r}_{0}\|_{\mathbf{H}}=\|\mathbf{r}_{i}\|_{\mathbf{H}}/\|\mathbf{b}\|_{\mathbf{H}}<10^{-10}$. The three considered meshes are the $32158$ vertex mesh that was used in the tests above as well as the two less refined meshes represented in Figure 3 (2373 and 8643 vertices). After elimination of the degrees of freedom on the boundary, the resulting linear systems have respectively 31502, 8307 and 2197 dofs. The problem is solved by WPD-GMRES without deflation ($\text{m}=0$ and with deflation of $\text{m}=100$ vectors). The results are presented in Table 5. We notice that the iteration counts increase weakly with the mesh size. For example, from 236 for 2,197 dofs to 275 for 32,502 dofs in the case where 100 vectors are deflated. We also report on the quantities that appear in the convergence bounds: $\kappa(\mathbf{H}_{\text{DD}}\mathbf{M})$, $|\lambda_{1}|=\rho(\mathbf{M}^{-1}\mathbf{N})$ and $|\lambda_{101}|$. We know from the theory in [28] that $\kappa(\mathbf{H}_{\text{DD}}\mathbf{M})$ is bounded independently of the mesh size and we indeed notice that it does not depend very much on the mesh. In [26, Section 5], it was proved that for this particular PDE, $|\lambda_{1}|=\rho(\mathbf{M}^{-1}\mathbf{N})$ is also bounded independently of the mesh size by
@@
Table 5: For $\eta=100$ and $\mathbf{H}=\mathbf{H}_{\text{DD}}$ in Section 7.3 , three mesh sizes are considered.
@@
Table 6: Iteration counts for weighted and unweighted algorithms. Both algorithms use the stopping criterion coming from the unweighted algorithm.
Figure 9: Comparison of weighted and unweighted algorithms, without deflation and with deflation of $100$ vectors. The chosen norm is the one that is minimized by the unweighted algorithm.